# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v1

@@ Appendix D Further Details: Network Architectures and Training
Autoencoder networks are studied in these results, due to their lack of one-hot classification head, which could impose anisotropy on results [^6] Some anisotropy may remain due to the spatialised nature of the image data; however, this is inherent to the data and the nature of visual datasets. Residual anisotropy through data is not the remit of the hypothesis being tested nor the overarching algebraic symmetry framework being studied. Though its effect was considered and discussed in App. A.. The autoencoders are comprised of an encoding, decoding, and hidden section. Separately categorising the ‘hidden’ layer is perhaps unorthodox but important for the study. All networks tested are derived from the one shown in Fig. 31. This diagrammatic depiction is consistent with open-sourced standardisation available [4] at https://github.com/GeorgeBird1/Diagramatic-Neural-Networks.
@@
Figure 31: Shows an autoencoding architecture for reconstruction of CIFAR data, in a standardised diagrammatic key system [4]. It is comprised of three blocks: encoder, hidden layers and decoder. The encoder flattens the tenser representations into a vector of dimension 3072 and then performs an affine transformation down to n dimensions. Following this, each hidden layer consists of the application of an activation function, which is either standard or isotropic-tanh, followed by another affine transformation from n to n dimensions. This hidden layer process is repeated $L$ times. Finally, the decoder applies the chosen activation function again, followed by an affine map from n to 3072 dimensions, before reshaping the vector into its classical representation. Various latent layers are drawn from the model, particularly after the encoder, denoted as layer 00, or after a hidden-layer block, denoted as $k$. Therefore, all latent layers of the study are $n$-dimensional. The parameters of all affine maps are optimisable through gradient descent.
This depicts the architectures used across all results, where $n$ varies in width and $L$ varies in depth of the autoencoder model. The latent layer of study is indexed by integers $k\in\left[0,L\right]$. For MNIST [11] models, the tensor representations are changed from $\left[32,32,1\right]\rightarrow\left[28,28,1\right]$ and $\left[3072\right]\rightarrow\left[784\right]$ in all instances, with corresponding changes in affine maps. This ensures these results follow from a similar and comparable model.
@@
Finally, in accordance with App. A, the MNIST and CIFAR datasets are sometimes normalised over their training sets to reduce statistical geometry artefacts. This involves an element-wise calculation of the mean tensor over all training samples, as well as for the standard deviation statistic. Where the standard deviation is found to be 00, those elements are replaced with the value $1$. When normalisation is used, every ingoing sample per batch is standardised by subtracting the mean and dividing by the standard deviation, where the non-zero standard deviation is made equal to one.
@@
Throughout all experiments, a batch size of $24$ was maintained, alongside a learning rate for PyTorch’s [12] momentum gradient descent, with momentum $0.9$, a learning rate $0.08$, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal [13] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances [2], although standard multivariate normal or orthogonal initialisation would have been equally valid choices, but Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions [2]. All values were kept in float64.
Networks were trained for a total of $125$ epochs; however, intermediate measurements were gathered after training for each of the incremental following amount of epochs: $\left[[1,1,2,2,2,5,5,6,6,10,10,10,15,15,15,20]\right]$ corresponding to a cumulative total of $\left[1,2,4,6,8,13,18,24,30,40,50,60,75,90,105,125\right]$. This was chosen arbitrarily, but was found to give meaningful insight into the formation of the representation structures.
@@
Data was gathered using combination-PPP, with $\epsilon=0.75$, which was found to be a good overall value. Plots were compiled using Gaussians centred at each representation’s projected coordinates, and the colour map indicates representation density per region, with a maximum value of the mean plus five standard deviations and a minimum of zero. Additionally, in the anisotropic cases, so many representations became aligned with the axes that computing the Gaussian maps became intractable due to the number of computations. Therefore, a random subset of the samples generated by PPP was plotted. The size of this subset was chosen to be the closest integer of the larger set raised to the power of $0.9$. Hence, the number of representations plotted monotonically increased with the number of PPP samples, yet remained computationally tractable for very large numbers of representations. This approach was used in all cases.

### v2

@@ Appendix D Further Details: Network Architectures and Training
Autoencoder networks are studied in these results, due to their lack of one-hot classification head, which could impose anisotropy on results [^9] Some anisotropy may remain due to the spatialised nature of the image data; however, this is inherent to the data and the nature of visual datasets. Residual anisotropy through data is not the remit of the hypothesis being tested nor the overarching algebraic symmetry framework being studied. Though its effect was considered and discussed in App. A.. The autoencoders are comprised of an encoding, decoding, and hidden section. Separately categorising the ‘hidden’ layer is perhaps unorthodox but important for the study. All networks tested are derived from the one shown in Fig. 31. This diagrammatic depiction is consistent with open-sourced standardisation available Bird [2025b] at https://github.com/GeorgeBird1/Diagramatic-Neural-Networks.
@@
Figure 31: Shows an autoencoding architecture for reconstruction of CIFAR data, in a standardised diagrammatic key system Bird [2025b]. It is comprised of three blocks: encoder, hidden layers and decoder. The encoder flattens the tensor representations into a vector of dimension 3072 and then performs an affine transformation down to n dimensions. Following this, each hidden layer consists of the application of an activation function, which is either standard or isotropic-tanh, followed by another affine transformation from n to n dimensions. This hidden layer process is repeated $L$ times. Finally, the decoder applies the chosen activation function again, followed by an affine map from n to 3072 dimensions, before reshaping the vector into its classical representation. Various latent layers are drawn from the model, particularly after the encoder, denoted as layer 0, or after a hidden-layer block, denoted as $k$. Therefore, all latent layers of the study are $n$-dimensional. The parameters of all affine maps are optimisable through gradient descent.
[…]
@@
Finally, following from App. A, the MNIST and CIFAR datasets are sometimes normalised over their training sets to reduce statistical geometry artefacts. This involves an element-wise calculation of the mean tensor over all training samples, as well as for the standard deviation statistic. Where the standard deviation is found to be 0, those elements are replaced with the value 111. When normalisation is used, every ingoing sample per batch is standardised by subtracting the mean and dividing by the standard deviation, where the non-zero standard deviation is made equal to one.
@@
Throughout all experiments, a batch size of 242424 was maintained, alongside a learning rate for PyTorch’s Paszke et al. [2019] momentum gradient descent, with momentum 0.90.90.9, a learning rate 0.080.080.08, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal Glorot and Bengio [2010] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances Bird [2025a]. However, standard multivariate normal or orthogonal initialisation would have been equally valid choices. Still, Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions Bird [2025a]. All values were kept in float64.
[…]
@@
Data was gathered using combination-PPP, with $\epsilon=0.75$, which was found to be a good overall value. Plots were compiled using Gaussians centred at each representation’s projected coordinates, and the colour map indicates representation density per region, with a maximum value of the mean plus five standard deviations and a minimum of zero. Additionally, in the anisotropic cases, so many representations became aligned with the axes that computing the Gaussian maps became intractable due to the number of computations. Therefore, a random subset of the samples generated by PPP was plotted. The size of this subset was chosen to be the closest integer of the larger set raised to the power of 0.90.90.9. Hence, the number of representations plotted monotonically increased with the number of PPP samples, yet remained computationally tractable for very large numbers of representations. This approach was used in all cases.